A teacher stands well back from an outside doorwayA teacher stands well back from an outside doorway 0.85 m wide, and blows a whistle of frequency 950?Hz. Ignoring reflections, estimate at what angle(s) ?mmin (0?<?mmin<90?) it is not possible to hear the whistle clearly on the playground outside the doorway. Assume 340 m/s for the speed of sound.
Answer #1

Solution:

The given case involves the diffraction of the sound waves by the doorway of width a. We need to find the minima of sound wave intensity at certain angle(s) θm,min where m is the order of diffraction.

Width of the doorway, a = 0.85 m

Frequency of the sound waves, f = 950 Hz = 950/s

Speed of sound waves, v = 340 m/s

Thus the wavelength λ of sound waves is given by,

v = f*λ

λ = v/f

λ = (340 m/s)/(950/s)

λ = 0.357895 m

The mth order minima occur at an angle of θm,min

a*sin(θm,min) = m*λ

sin(θm,min) = m*λ/a

Now for first order diffraction, m = 1, we have

sin(θ1,min) = (1)*( 0.357895 m) / (0.85 m)

sin(θ1,min) = 0.421053

θ1,min = sin-1(0.421053)

θ1,min = 24.9011o

With two significant digits,

θ1,min = 25o

For second order diffraction, m = 2, we have

sin(θ2,min) = (2)*( 0.357895 m) / (0.85 m)

sin(θ2,min) = 0.842106

θ2,min = sin-1(0.842106)

θ2,min = 57.3632o

With two significant digits,

θ2,min = 57o                         

For m = 3

sin(θ3,min) = (3)*( 0.357895 m) / (0.85 m)

sin(θ3,min) = 1.26316

One cannot take sin inverse of a number greater than 1 (or less than -1), it tells us that only two minima occur at either sides of doorway at,

θ1,min = 25o   and θ2,min = 57o

Hence the answers are

25o, 57o